OpenStudy (anonymous):

Which equation would you use to solve the following word problem? Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $22.95 per month plus an additional $0.08 per minute. Cell Best has a plan that costs $27.95 per month plus an additional $0.06 per minute. How many minutes can Sasha talk and have the same cost each month?

OpenStudy (anonymous):

Where are the equations?

OpenStudy (anonymous):

There Not One

OpenStudy (anonymous):

"Which equation would you use to solve the following word problem?" Are you sure?

OpenStudy (anonymous):

These Are The Answer 0.14m = 5.00 22.95 + 0.08m = 27.95 + 0.06m 22.95 + 0.08m = 28.01 27.95 + 0.06m = 23.03

OpenStudy (anonymous):

The answer is the second one. 22.95 + 0.08m = 27.95 + 0.06m

OpenStudy (anonymous):

OK Thank You!

OpenStudy (anonymous):

No problem mate.